# Accelerating Equity: Overcoming the Gender Gap in VC Funding

### v1

@@ 6.1 Mechanism
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Accelerators with larger cohorts More experienced accelerators Funding Within One Year Funding Within Five Years Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) Exceeds (0/1) Exceeds (0/1) Two Mil. Five Mil. Two Mil. Five Mil. Two Mil. Five Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.072* -0.044 -0.044 -0.110 -0.093** -0.036 0.000 0.125 (0.043) (0.033) (0.149) (0.103) (0.043) (0.031) (0.303) (0.297) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.018 0.037 0.069 0.053 -0.063 -0.041 0.262 0.342 (0.055) (0.039) (0.217) (0.162) (0.126) (0.081) (0.429) (0.406) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.053 -0.006 -0.025 -0.056 -0.156 -0.076 0.262 0.466 (0.062) (0.050) (0.196) (0.137) (0.141) (0.096) (0.515) (0.493) Sample Large cohorts (above median) Experienced accelerator (above median) $N$ 408 408 408 408 326 326 326 326 $R^{2}$ 0.082 0.061 0.060 0.071 0.076 0.060 0.098 0.088 * – $p<0.1$; ** – $p<0.05$; *** – $p<0.01$
  • Notes: This table presents linear probability estimates from (3) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Columns 1 to 4 focus on startups in accelerators with cohorts above the median cohort size. Columns 5 to 8 focus on startups in accelerators with more than the median number of years of activity. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
@@ 6.2 Economic Benefits of Bridging the Gender Gap
The remaining columns further investigate these findings by conditioning on surviving startups. Columns (4) to (6) confirm that female-founded startups receive less funding, particularly struggling to reach the five-million milestone. Despite this funding gap, female-founded startups are just as likely to survive or successfully exit within five years as other startups. This suggests that female-founded startups can achieve similar outcomes as male-founded or mixed-gender startups with fewer financial resources, highlighting their high productivity.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Failed in its First Acquired Funding Exceeds (0/1) Failed Exited 1 Year 5 Years in 5 Years One Million Two Millions Five Millions (0/1) (0/1) Female Founder 0.068 0.029 0.068 -0.185 -0.133 -0.213* -0.026 -0.015 (0.128) (0.137) (0.134) (0.186) (0.184) (0.129) (0.141) (0.162) Other Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Sample All All All Conditional on surviving in the first year since graduation $N$ 736 736 736 325 325 325 325 325 $R^{2}$ 0.044 0.108 0.044 0.091 0.061 0.091 0.072 0.055 * – $p<0.1$; ** – $p<0.05$; *** – $p<0.01$
  • Notes: This table presents linear probability estimates from (3) of startup outcomes, including failure, acquisition, and funding milestones. The first three columns analyze failure within one and five years, as well as acquisition within five years. Columns (4)–(6) examine whether a startup raised more than \$1 million, \$2 million, or \$5 million, conditional on surviving the first year. The last two columns report failure and exit probabilities. Bootstrapped standard errors are in parentheses. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors in parentheses.
@@ 7 Conclusions
The methodology used in this paper combines a one-to-many matching model with non-transferable utility and a control function approach to estimate startup outcomes. By first modeling the matching process between startups and accelerators, and then using this information to control for unobserved startup quality, we provide a robust analysis of funding disparities. This approach allows us to account for the endogenous factors that may affect venture outcomes and offers a nuanced understanding of the gender funding gap in the venture capital landscape.
@@ A.3 Omitted Tables
(1) (2) (3) (4)
Acc in Democratic State 0.007 0.012 0.025 0.028
(0.032) (0.032) (0.032) (0.033)
Year Fixed Effects Y Y
$N$ 736 736 74 74
$R^{2}$ 0.000 0.002 0.009 0.022
* – $p<0.1$; ** – $p<0.05$; *** – $p<0.01$
  • Notes: This table shows the regression coefficients of voting for the democratic party in the 2008 presidential election. The dependent variables for the first two models are indicators of whether the accelerators’ participating startups are founded by women. The dependent variables for the last two models are the percentages of women-founded startups.
@@
Women Men p-value
Founded Founded of W>M
log(InvestSize 1yr)
Mean 6.427 6.562 0.293
Variance 0.777 1.624 0.027
Skewness -1.065 -0.397 0.114
log(InvestSize 5yr)
Mean 7.388 7.768 0.160
Variance 1.602 3.768 0.004
Skewness -0.204 0.048 0.032
  • Notes: This table shows the log(InvestSize) distribution moments for VC funded startups. The last column presents the p value of whether the corresponding moment for women-founded startups is larger than that of their male counterparts.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6)
Funding Within One Year Funding Within Five Years
Exceeds (0/1) Exceeds (0/1)
One Mil. Two Mil. Five Mil. One Mil. Two Mil. Five Mil.
Female Founder ($\alpha_{1}$) -0.125 -0.130* -0.065 -0.174 -0.044 -0.206
(0.182) (0.076) (0.053) (0.202) (0.198) (0.165)
Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.113 -0.022 0.041 -0.026 -0.214 -0.018
(0.202) (0.086) (0.048) (0.206) (0.230) (0.188)
Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$
Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Test
$H_{0}:\alpha_{1}+\alpha_{2}=0$0.012 -0.152 -0.024 -0.199 -0.257 -0.223
(0.220) (0.088) (0.066) (0.215) (0.207) (0.159)
Sample Conditional on surviving in the first year since graduation
$N$ 325 325 325 325 325 325
$R^{2}$ 0.161 0.103 0.062 0.091 0.064 0.091
* – $p<0.1$; ** – $p<0.05$; *** – $p<0.01$
  • Notes: This table presents linear probability estimates from (3) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Only startup that survived the first year after graduation are considered. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
@@
Proof. Observe that through law of iterated expectation, we have
@@
The second equality is due to the fact that the $\sigma$-field generated by $(\epsilon,X)$ determines $\mu$. The third equality is due to the fact that $(\epsilon^{a^{\prime}s^{\prime}},X)$ are independent of $\eta^{as}$ when $a^{\prime}s^{\prime}\neq as$. The fourth equality follows from properties of bivariate normal distribution. Consider the product $\rho\sigma$ as a deterministic parameter different from zero; this completes the proof. $\Box$
@@ A.6 The Matching Model Estimates
Accelerator w Female Founder*Female Founder -0.006 0.165

### v2

@@ 6.1 Mechanism
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Accelerators with larger cohorts More experienced accelerators Funding Within One Year Funding Within Five Years Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) Exceeds (0/1) Exceeds (0/1) Two Mil. Five Mil. Two Mil. Five Mil. Two Mil. Five Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.072 -0.044 -0.044 -0.110 -0.093 -0.036 0.000 0.125 (0.043) (0.033) (0.149) (0.103) (0.043) (0.031) (0.303) (0.297) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.018 0.037 0.069 0.053 -0.063 -0.041 0.262 0.342 (0.055) (0.039) (0.217) (0.162) (0.126) (0.081) (0.429) (0.406) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.053 -0.006 -0.025 -0.056 -0.156 -0.076 0.262 0.466 (0.062) (0.050) (0.196) (0.137) (0.141) (0.096) (0.515) (0.493) Sample Large cohorts (above median) Experienced accelerator (above median) $N$ 408 408 408 408 326 326 326 326 $R^{2}$ 0.082 0.061 0.060 0.071 0.076 0.060 0.098 0.088 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Columns 1 to 4 focus on startups in accelerators with cohorts above the median cohort size. Columns 5 to 8 focus on startups in accelerators with more than the median number of years of activity. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
@@ 6.2 Economic Benefits of Bridging the Gender Gap
The remaining columns further investigate these findings by conditioning on surviving startups. Columns (4) to (6) confirm that female-founded startups receive less funding, particularly struggling to reach the five-million milestone. Despite this funding gap, female-founded startups are just as likely to survive or successfully exit within five years as other startups. This suggests that female-founded startups can achieve similar outcomes as male-founded startups with fewer financial resources, highlighting their high productivity.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Failed in its First Acquired Funding Exceeds (0/1) Failed Exited 1 Year 5 Years in 5 Years One Million Two Millions Five Millions (0/1) (0/1) Female Founder 0.068 0.029 0.068 -0.185 -0.133 -0.213 -0.026 -0.015 (0.128) (0.137) (0.134) (0.186) (0.184) (0.129) (0.141) (0.162) Other Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Sample All All All Conditional on surviving in the first year since graduation $N$ 736 736 736 325 325 325 325 325 $R^{2}$ 0.044 0.108 0.044 0.091 0.061 0.091 0.072 0.055 Notes: This table presents linear probability estimates from (4) of startup outcomes, including failure, acquisition, and funding milestones. The first three columns analyze failure within one and five years, as well as acquisition within five years. Columns (4)–(6) examine whether a startup raised more than \$1 million, \$2 million, or \$5 million, conditional on surviving the first year. The last two columns report failure and exit probabilities. Bootstrapped standard errors are in parentheses. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors in parentheses.
@@ 7 Conclusions
We also contributed by offering a computationally light approach to estimating matching designs in large markets. In particular, the methodology used in this paper combines a one-to-many matching model with non-transferable utility and a control function approach to estimate startup outcomes. By first modeling the matching process between startups and accelerators, and then using this information to control for unobserved startup quality, we provide a robust analysis of funding disparities. This approach allows us to account for the endogenous factors that may affect venture outcomes and offers a nuanced understanding of the gender funding gap in the venture capital landscape.
@@ A.3 Omitted Tables
Indicator for Share of Female Founded Startups Female Founded Sturtups (1) (2) (3) (4) Acc in Democratic State 0.007 0.012 0.025 0.028 (0.032) (0.032) (0.032) (0.033) Year Fixed Effects Y Y $N$ 736 736 74 74 $R^{2}$ 0.000 0.002 0.009 0.022 Notes: This table shows the coefficients from regressing the occurrence of female founded startups on a dummy for voting for the democratic party in the 2008 presidential election. The dependent variables for the first two models are indicators of whether the accelerators’ participating startups are founded by women. The dependent variables for the last two models are the percentages of women-founded startups. Even columns include state-fixed effects.
@@
Women Men p-value Founded Founded of W¿M log(InvestSize 1yr) Mean 6.427 6.562 0.293 Variance 0.777 1.624 0.027 Skewness -1.065 -0.397 0.114 log(InvestSize 5yr) Mean 7.388 7.768 0.160 Variance 1.602 3.768 0.004 Skewness -0.204 0.048 0.032 Notes: This table shows the log(InvestSize) distribution moments for VC funded startups. The last column presents the p value of whether the corresponding moment for women-founded startups is larger than that of their male counterparts.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6) Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) One Mil. Two Mil. Five Mil. One Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.125 -0.130 -0.065 -0.174 -0.044 -0.206 (0.182) (0.076) (0.053) (0.202) (0.198) (0.165) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.113 -0.022 0.041 -0.026 -0.214 -0.018 (0.202) (0.086) (0.048) (0.206) (0.230) (0.188) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.012 -0.152 -0.024 -0.199 -0.257 -0.223 (0.220) (0.088) (0.066) (0.215) (0.207) (0.159) Sample Conditional on surviving in the first year since graduation $N$ 325 325 325 325 325 325 $R^{2}$ 0.161 0.103 0.062 0.091 0.064 0.091 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Only startup that survived the first year after graduation are considered. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
@@
Proof.
Observe that through law of iterated expectation, we have
@@
The second equality is due to the fact that the $\sigma$-field generated by $(\epsilon,X)$ determines $\mu$. The third equality is due to the fact that $(\epsilon^{a^{\prime}s^{\prime}},X)$ are independent of $\eta^{as}$ when $a^{\prime}s^{\prime}\neq as$. The fourth equality follows from properties of bivariate normal distribution. Consider the product $\rho\sigma$ as a deterministic parameter different from zero; this completes the proof. ∎
@@ A.6 The Matching Model Estimates
Accelerator w Female Founder×$\times$×Female Founder -0.006 0.165